A formula for morphine sulfate and ibuprofen injection for patient-controlled analgesia is as follows:

Morphine sulfate 200 mg
Ibuprofen 8 mg
Lactic acid qs
0.9% sodium chloride injection qs 100 mL

A patient has been prescribed 2 mg of morphine sulfate every 2 hours as needed for severe pain. The purpose of the lactic acid in this formulation is to adjust the pH to the range of 3 to 3.8, and lactic acid is supplied as an 85% w/w solution with a specific gravity of approximately 1.20. If 0.55 mL of lactic acid solution is used to adjust the pH of the solution, what is the final percent w/v strength of lactic acid in the mixture?
Answer #1

Answer ,The volume of 85%w/w lactic acid solution required toh adjust the pH is ,as given = 0.55 ml

Let assume the measured density of lactic acid solution be 1 g/ml.

Thus , the weight of 0.55 ml solution of lactic acid is = density × volume = 1× 0.55 = 0.55 g

As given in question, the concentration of la tic acid is 85% w/w the amount of lactic acid present in 0.55 g of lactic acid solution is given by , Weight of lactic acid = (85 × 0.55)/100 = 0.4675 g

Thus the percent weight per volume of final solution is given by ,

(0.4675 /1 ) × 100 = 46.75 %

Ansser:- the concentration of lactic acid in the prepared final solution was found to be 46.75%